OpenStudy (anonymous):

I'm having a little bit of trouble with logs (Well, more specifically, rewriting logs in terms of others). log 6 (10)=1.3 log 6 (7)=1.1 log6 (8)=1.2 Find log 6 (3/50)

OpenStudy (anonymous):

the gimmick is to figure out how two write \(\frac{3}{50}\) in terms of \(10, 7,8\)

OpenStudy (anonymous):

oh and i guess \(6\) since that is the base

OpenStudy (anonymous):

one question is it asking this ? \[\log_{6} (3/5)\]

OpenStudy (anonymous):

or is the 6 multiplies by the (3/5)

OpenStudy (anonymous):

gimmick is to do this write \(\frac{3}{50}=\frac{6}{100}=\frac{6}{10^2}\)

OpenStudy (anonymous):

The first one you said. except it's log6 (3/50)

OpenStudy (anonymous):

@satellite73: Alright :) What next?

OpenStudy (anonymous):

and so \[\log_6(\frac{3}{50})=\log_6(\frac{6}{10^2})\] now break it apart

OpenStudy (anonymous):

you get \[\log_6(\frac{6}{10^2})=\log_6(6)-\log_6(10^2)=\log_6(6)-2\log(10)\]

OpenStudy (anonymous):

you have all the numbers you need, since you are told \(\log_6(10)=1.3\) ( a lie, but no matter) and \(\log_6(6)\) should be obvious

OpenStudy (anonymous):

Ah, okay, got it! I have one other question I don't get. log8 (12) = 1.2 log8 (7) = 0.9 log8 (9) = 1.1 Find log8 (81/7)